# Radio Map-Assisted CSI Tracking with Uncertain Locations in Massive MIMO Networks

### v1

@@ Appendix C Proof of Proposition 2
Recall that ${\bf D}_{k_{i}}={\bf A}{\bf C}_{k_{i}}{\bf A}^{\text{H}}+\sigma^{2}{\bf I}_{M\times M}$ and ${\bf C}_{k_{i}}={\bf U}_{k_{i}}{\bf\Sigma}_{k_{i}}{\bf U}_{k_{i}}^{\text{H}}$. Define ${\bf V}_{k_{i}}\triangleq{\bf A}{\bf U}_{k_{i}}{\bf\Sigma}_{k_{i}}^{1/2}$, and thus, ${\bf D}_{k_{i}}$ is rewritten as
${\bf D}_{k_{i}}={\bf V}_{k_{i}}{\bf V}_{k_{i}}^{\text{H}}+\sigma^{2}{\bf I}_{M\times M}.$ (45)
Setting the derivative of $F(\mathcal{C}_{R})$ with respect to (w.r.t.) ${\bf V}_{k_{i}}$ to 00, one can obtain
$\frac{\partial F}{\partial{\bf V}_{k_{i}}}={\bf D}_{k_{i}}^{-1}{\bf M}_{k_{i}}{\bf D}_{k_{i}}^{-1}{\bf V}_{k_{i}}-{\bf D}_{k_{i}}^{-1}{\bf V}_{k_{i}}=0$ (46)
where ${\bf M}_{k_{i}}$ is the average sampling covariance defined in (28).
Hence, the critical point of $F(\mathcal{C}_{T})$ w.r.t. ${\bf V}_{k_{i}}$ satisfies
${\bf M}_{k_{i}}{\bf D}_{k_{i}}^{-1}{\bf V}_{k_{i}}={\bf V}_{k_{i}}.$ (47)
Since $({\bf A}{\bf U}_{k_{i}})({\bf A}{\bf U}_{k_{i}})^{\text{H}}={\bf A}{\bf U}_{k_{i}}{\bf U}_{k_{i}}^{\text{H}}{\bf A}={\bf I}_{M\times M}$, it follows that ${\bf A}{\bf U}_{k_{i}}$ is semi-unitary. Therefore, we have
${\bf D}_{k_{i}}^{-1}{\bf V}_{k_{i}}$ $=\left({\bf V}_{k_{i}}{\bf V}_{k_{i}}^{\text{H}}+\sigma^{2}{\bf I}_{M\times M}\right)^{-1}{\bf V}_{k_{i}}$
$=\left({\bf A}{\bf U}_{k_{i}}\left({\bf\Sigma}_{k_{i}}+\sigma^{2}{\bf I}_{N_{t}\times N_{t}}\right){\bf U}_{k_{i}}^{\text{H}}{\bf A}^{\text{H}}\right)^{-1}{\bf V}_{k_{i}}$
$=\left({\bf A}{\bf U}_{k_{i}}\right)\left({\bf\Sigma}_{k_{i}}+\sigma^{2}{\bf I}_{N_{t}\times N_{t}}\right)^{-1}\left({\bf A}{\bf U}_{k_{i}}\right)^{\text{H}}{\bf V}_{k_{i}}$
$={\bf A}{\bf U}_{k_{i}}\left({\bf\Sigma}_{k_{i}}+\sigma^{2}{\bf I}_{N_{t}\times N_{t}}\right)^{-1}{\bf U}_{k_{i}}^{\text{H}}{\bf A}^{\text{H}}{\bf A}{\bf U}_{k_{i}}{\bf\Sigma}_{k_{i}}^{1/2}$ (48)
$={\bf V}_{k_{i}}({\bf\Sigma}_{k_{i}}+\sigma^{2}{\bf I}_{N_{t}\times N_{t}})^{-1}{\bf U}_{k_{i}}^{\text{H}}{\bf A}^{\text{H}}{\bf A}{\bf U}_{k_{i}}$ (49)
@@
${\bf P}_{M}={\bf U}_{k_{i}}^{\text{H}}{\bf A}^{\text{H}}{\bf A}{\bf U}_{k_{i}}=\left[\begin{array}[]{cc}{\bf I}_{M\times M}&{\bf 0}\\ {\bf 0}&{\bf 0}\end{array}\right].$ (50)
Substituting (50) into (48), it follows that ${\bf D}_{k_{i}}^{-1}{\bf V}_{k_{i}}={\bf V}_{k_{i}}({\bf\Sigma}_{k_{i}}+\sigma^{2}{\bf I}_{N_{t}\times N_{t}})^{-1}{\bf P}_{M}$, and thus, (47) can be simplified as
${\bf M}_{k_{i}}({\bf A}{\bf U}_{k_{i}})=({\bf A}{\bf U}_{k_{i}})({\bf\Sigma}_{k_{i}}{\bf P}_{M}+\sigma^{2}{\bf P}_{M}).$ (51)
@@
  • [9] J. Vinogradova, G. Fodor, and P. Hammarberg, “On estimating the autoregressive coefficients of time-varying fading channels,” pp. 1–5, Helsinki, Finland, 2022.
@@
  • [12] T. Peng, R. Zhang, X. Cheng, and L. Yang, “Lstm-based channel prediction for secure massive mimo communications under imperfect csi,” pp. 1–6, Dublin, Ireland, 2020.
@@
  • [28] T. Li, X. Wang, P. Fan, and T. Riihonen, “Position-aided channel estimation for large-scale MIMO in high-speed railway scenarios,” pp. 1–6, Washington, DC, USA, 2016.
@@
  • [30] S. Taner, M. Guillaud, O. Tirkkonen, and C. Studer, “Channel charting for streaming CSI data,” pp. 1648–1653, 2023.
  • [31] M. Stahlke, G. Yammine, T. Feigl, B. M. Eskofier, and C. Mutschler, “Velocity-based channel charting with spatial distribution map matching,” IEEE J Indoor Seamless Position Navig., vol. 2, pp. 230–239, 2024.
  • [32] S. Taner, V. Palhares, and C. Studer, “Channel charting in real-world coordinates,” pp. 3940–3946, 2023.
@@
  • [36] T. Choi, I. Kanno, M. Ito, W.-Y. Chen, and A. F. Molisch, “A realistic path loss model for cell-free massive MIMO in urban environments,” IEEE Global Commun. Conf., pp. 2468–2473, 2022.

### v2

@@ Appendix C Proof of Proposition 2
Recall that ${\bf D}_{i}^{(q)}={\bf A}{\bf C}_{i}^{(q)}{\bf A}^{\text{H}}+\sigma^{2}{\bf I}_{M\times M}$ and ${\bf C}_{i}^{(q)}={\bf U}_{i}^{(q)}\bm{\Sigma}_{i}^{(q)}({\bf U}_{i}^{(q)})^{\text{H}}$. Define ${\bf V}_{i}^{(q)}\triangleq{\bf A}{\bf U}_{i}^{(q)}(\bm{\Sigma}_{i}^{(q)})^{1/2}$, and thus, ${\bf D}_{i}^{(q)}$ is rewritten as
${\bf D}_{i}^{(q)}={\bf V}_{i}^{(q)}({\bf V}_{i}^{(q)})^{\text{H}}+\sigma^{2}{\bf I}_{M\times M}.$ (46)
Setting the derivative of $F(\{{\bf C}_{i}^{(q)}\}_{i\in\mathcal{N}})$ with respect to (w.r.t.) ${\bf V}_{i}^{(q)}$ to 00, one can obtain
$\frac{\partial F}{\partial{\bf V}_{i}^{(q)}}=({\bf D}_{i}^{(q)})^{-1}{\bf M}_{i}^{(q)}({\bf D}_{i}^{(q)})^{-1}{\bf V}_{i}^{(q)}-({\bf D}_{i}^{(q)})^{-1}{\bf V}_{i}^{(q)}=0$ (47)
where ${\bf M}_{i}^{(q)}$ is the average sampling covariance defined in (28).
Hence, the critical point of $F(\{{\bf C}_{i}^{(q)}\}_{i\in\mathcal{N}})$ w.r.t. ${\bf V}_{i}^{(q)}$ satisfies
${\bf M}_{i}^{(q)}({\bf D}_{i}^{(q)})^{-1}{\bf V}_{i}^{(q)}={\bf V}_{i}^{(q)}.$ (48)
Since $({\bf A}{\bf U}_{i}^{(q)})({\bf A}{\bf U}_{i}^{(q)})^{\text{H}}={\bf A}{\bf U}_{i}^{(q)}({\bf U}_{i}^{(q)})^{\text{H}}{\bf A}^{\text{H}}={\bf I}_{M\times M}$, it follows that ${\bf A}{\bf U}_{i}^{(q)}$ is semi-unitary. Therefore, we have
$({\bf D}_{i}^{(q)})^{-1}{\bf V}_{i}^{(q)}$ $=\left({\bf V}_{i}^{(q)}({\bf V}_{i}^{(q)})^{\text{H}}+\sigma^{2}{\bf I}_{M\times M}\right)^{-1}{\bf V}_{i}^{(q)}$
$=\left({\bf A}{\bf U}_{i}^{(q)}\tilde{\boldsymbol{\Sigma}}_{i}^{(q)}({\bf U}_{i}^{(q)})^{\text{H}}{\bf A}^{\text{H}}\right)^{-1}{\bf V}_{i}^{(q)}$
$=\left({\bf A}{\bf U}_{i}^{(q)}\right)\left(\tilde{\boldsymbol{\Sigma}}_{i}^{(q)}\right)^{-1}\left({\bf A}{\bf U}_{i}^{(q)}\right)^{\text{H}}{\bf V}_{i}^{(q)}$
$={\bf A}{\bf U}_{i}^{(q)}\left(\tilde{\boldsymbol{\Sigma}}_{i}^{(q)}\right)^{-1}{\bf P}_{M}(\bm{\Sigma}_{i}^{(q)})^{1/2}$ (49)
$={\bf V}_{i}^{(q)}(\tilde{\boldsymbol{\Sigma}}_{i}^{(q)})^{-1}{\bf P}_{M}$ (50)
where $\tilde{\boldsymbol{\Sigma}}_{i}^{(q)}=\bm{\Sigma}_{i}^{(q)}+\sigma^{2}{\bf I}_{N_{t}\times N_{t}}$ and ${\bf P}_{M}=({\bf U}_{i}^{(q)})^{\text{H}}{\bf A}^{\text{H}}{\bf A}{\bf U}_{i}^{(q)}$.
@@
${\bf P}_{M}=({\bf U}_{i}^{(q)})^{\text{H}}{\bf A}^{\text{H}}{\bf A}{\bf U}_{i}^{(q)}=\left[\begin{array}[]{cc}{\bf I}_{M\times M}&{\bf 0}\\ {\bf 0}&{\bf 0}\end{array}\right].$ (51)
Substituting (51) into (49), it follows that $({\bf D}_{i}^{(q)})^{-1}{\bf V}_{i}^{(q)}={\bf V}_{i}^{(q)}(\bm{\Sigma}_{i}^{(q)}+\sigma^{2}{\bf I}_{N_{t}\times N_{t}})^{-1}{\bf P}_{M}$, and thus, (48) can be simplified as
${\bf M}_{i}^{(q)}({\bf A}{\bf U}_{i}^{(q)})=({\bf A}{\bf U}_{i}^{(q)})({\bf\Sigma}_{i}^{(q)}{\bf P}_{M}+\sigma^{2}{\bf P}_{M}).$ (52)
@@
  • [9] J. Vinogradova, G. Fodor, and P. Hammarberg, “On estimating the autoregressive coefficients of time-varying fading channels,” in Proc. IEEE Veh. Technol. Conf. (VTC), pp. 1–5, Helsinki, Finland, 2022.
@@
  • [12] T. Peng, R. Zhang, X. Cheng, and L. Yang, “LSTM-based channel prediction for secure massive MIMO communications under imperfect CSI,” in Proc. IEEE Int. Conf. Commun. (ICC), pp. 1–6, Dublin, Ireland, 2020.
@@
  • [28] T. Li, X. Wang, P. Fan, and T. Riihonen, “Position-aided channel estimation for large-scale MIMO in high-speed railway scenarios,” in Proc. IEEE Global Commun. Conf. (GLOBECOM), pp. 1–6, Washington, DC, USA, 2016.
@@
  • [30] S. Taner, M. Guillaud, O. Tirkkonen, and C. Studer, “Channel charting for streaming CSI data,” in Proc. Asilomar Conf. Signals Syst Comput., pp. 1648–1653, 2023.
  • [31] M. Stahlke, G. Yammine, T. Feigl, B. M. Eskofier, and C. Mutschler, “Velocity-based channel charting with spatial distribution map matching,” IEEE J. Indoor Seamless Position Navig., vol. 2, pp. 230–239, 2024.
  • [32] S. Taner, V. Palhares, and C. Studer, “Channel charting in real-world coordinates,” in Proc. IEEE Global Commun. Conf., pp. 3940–3946, 2023.
@@
  • [36] T. Choi, I. Kanno, M. Ito, W.-Y. Chen, and A. F. Molisch, “A realistic path loss model for cell-free massive MIMO in urban environments,” in Proc. IEEE Global Commun. Conf. (GLOBECOM), pp. 2468–2473, 2022.